OpenStudy (anonymous):

Choose the correct solution in roster form. N is the set of real numbers that are factors of 20. (1 point){2, 4, 5, 10} {2, 4, 6, 8, 10, 12, 14, 16, 18, 20} {1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20} {1, 2, 4, 5, 10, 20}

OpenStudy (swissgirl):

Well lets decide what the factors of 20 are

OpenStudy (anonymous):

ummm wait its 1, 2, 4, 5, 10, 20

OpenStudy (swissgirl):

1*20=20 2*10=20 4*5=20 So the factors of 20 are {1,2,4,5,10,20}

OpenStudy (swissgirl):

Correct :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

It is Good Day

OpenStudy (swissgirl):

hahahahah Its a good day for you :P lol

OpenStudy (anonymous):

Suppose U = {1, 2, 3, 4, 5, 6, 7, 8} is the universal set, and P = {1, 3, 5, 7}. What is P'? (1 point){2, 4, 6, 8} {1, 2, 3, 4, 5, 6, 7, 8} {1, 3, 5, 7} {1, 3, 5, 7, 8}

OpenStudy (swissgirl):

Well U just told me what P is so why are you asking what P is?

OpenStudy (swissgirl):

So what do u think P is?

OpenStudy (anonymous):

1 3 5 7

OpenStudy (swissgirl):

BRAVVVOOOOO

OpenStudy (anonymous):

wow didnt expect to be fullly acurate on that sweet